# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v1

@@ F.1 Multi-Modal Task Categories
F.2 Details of Continual Pretrained Model Evaluation
F.2.1 Descriptions of Zero-shot Evaluation Tasks
@@
F.3 Descriptions of Few-shot Evaluation Tasks
@@
F.3.1 Prompts for Evaluation Datasets
For detailed prompts for evaluation datasets, please see Table 13 for 0-shot, and Table 14 for few-shots.
@@ F.3.3 Few-shots Performance
F.4 Prompts for Multimodal Benchmark Evaluation
@@
F.5 Prompts for Generating LLM Answers
@@ F.6 Example of LLM Answers
F.7 Prompts for Evaluating LLM Answers
@@
Appendix G Details for Trading in the FinMem task
@@
Data: We evaluated the model’s performance on the FinMem task using the following datasets:
@@
Figure 8 and 9 present our prompts for trading in the FinMem task.
Initialize Profile 1. Operations: - Provide a performance overview of the trading stock based on available data. - Set up the risk inclination as the key character of the trading agent. 2. Range: Financial information such as the financial sectors, historical performance, and previous stock trends of the trading stock. 3. Prompts: You are an experienced trading manager and investment firm. Your task is to make informed decisions on the given stock based on the provided information. Under Self-Adaptive Risk Character Setting: When historical momentum is positive, you are a risk-seeking investor. But when historical momentum is negative, you are a risk-averse investor. 4. General background setting: You have accumulated a lot of information about the following sectors, so you are especially good at trading them: 1)Electric Vehicles (Automotive Sector). 2) Energy Generation and Storage…From year 2021 to 2022 September, Tesla’s continued growth and solid financial performance over the defined period …
Summarize 1. Operations: - Summarize different types of input information. - Distribute them to corresponding layers of the long-term memory database. 2. Range: Daily market news, Long Documents such as company 10-K and 10-Q reports 3. Prompts:
@@ 4. Outputs:
Appendix H Trading comparison in the FinMem Task
Figure 10 through 12 illustrate the overall cumulative returns over time for the FinMem task using different models. For TSLA and NIO stocks, we can see that FinLLaMA consistently outperforms other models across all time periods. On the COIN stock, FinLLaMA exhibits a more stable and consistent upward trend in cumulative returns compared to other models, particularly before February 2023. In contrast, for GOOG stock, while FinLLaMA’s cumulative return is slightly lower than that of LLAMA3-8B and Palmyra-Fin-70B-32K across various time periods, it remains superior to the Buy & Hold strategy.

### v2

@@ F.3 Multimodal Task Categories
F.4 Details of Continual Pretrained Model Evaluation
F.4.1 Descriptions of Zero-shot Evaluation Tasks
@@
F.4.2 Descriptions of Few-shot Evaluation Tasks
@@
For detailed prompts for evaluation datasets, please see Table 13 for 0-shot, and Table 14 for few-shots. The results are listed in Figure 5 and Figure 6.
@@
F.5 Prompts for Multimodal Benchmark Evaluation
@@
F.6 Prompts for Generating LLM Answers
@@ F.7 Example of LLM Answers
F.8 Prompts for Evaluating LLM Answers
@@
Appendix G Details for Trading in the Decision Making Task
@@
Data: We evaluated the model’s performance on the Decision Making task using the following datasets:
@@
Figure 8 and 9 present our prompts for trading in the Decision Making task.
Initialize Profile 1. Operations: - Provide a performance overview of the trading stock based on available data. - Set up the risk inclination as the key character of the trading agent. 2. Range: Financial information such as the financial sectors, historical performance, and previous stock trends of the trading stock. 3. Prompts: You are an experienced trading manager and investment firm. Your task is to make informed decisions on the given stock based on the provided information. Under Self-Adaptive Risk Character Setting: When historical momentum is positive, you are a risk-seeking investor. But when historical momentum is negative, you are a risk-averse investor. 4. General background setting: You have accumulated a lot of information about the following sectors, so you are especially good at trading them: 1)Electric Vehicles (Automotive Sector). 2) Energy Generation and Storage…From year 2021 to 2022 September, Tesla’s continued growth and solid financial performance over the defined period … Summarize 1. Operations: - Summarize different types of input information. - Distribute them to corresponding layers of the long-term memory database. 2. Range: Daily market news, Long Documents such as company 10-K and 10-Q reports 3. Prompts:
@@ 4. Outputs:
Appendix H Trading Comparison in the Decision Making Task
Figure 10 through 12 illustrate the overall cumulative returns over time for the Decision Making task using different models. For TSLA and NIO stocks, we can see that FinLLaMA consistently outperforms other models across all time periods. On the COIN stock, FinLLaMA exhibits a more stable and consistent upward trend in cumulative returns compared to other models, particularly before February 2023. In contrast, for GOOG stock, while FinLLaMA’s cumulative return is slightly lower than that of LLAMA3-8B and Palmyra-Fin-70B-32K across various time periods, it remains superior to the Buy & Hold strategy.
@@ Appendix H Trading Comparison in the Decision Making Task
Appendix I Ethical Statement
The development and dissemination of the Open FinLLMs by the authors carry full responsibility for any potential violation of rights or arising legal issues. All raw data we used are publicly available and do not contain any personal information. Diligent efforts have been undertaken to ensure the construction of the Open FinLLMs respects privacy and conforms to established ethical guidelines. The datasets compiled within Open FinLLMs are shared under the MIT license, with the expectation that users agree to adhere to its conditions.
This manuscript, including any associated large language models, source codes, datasets, and appendices ("Material"), is intended solely for academic research and educational purposes. Users must recognize that the Material is not designed to provide financial, legal, medical, or other professional advice, nor should it serve as a substitute for expert judgment or decision-making in any domain.
While the authors have taken reasonable measures to ensure the accuracy and reliability of the Material, no guarantees—explicit or implied—are provided regarding its completeness, accuracy, or fitness for specific purposes. The authors and their affiliated institutions disclaim responsibility for any outcomes, whether direct or indirect, arising from the use, misuse, or reliance upon the Material. Users are strongly advised to consult relevant professionals when making decisions based on outputs generated by this Material.
By accessing or utilizing this Material, individuals agree to indemnify, defend, and hold harmless the authors, as well as their affiliated organizations, from any claims, liabilities, or damages resulting from the use or interpretation of the Material, including outputs produced by the large language model. Users bear the responsibility for ensuring the appropriate and ethical use of this Material in accordance with applicable laws and regulations.